# Multistability of graphene nanobubbles

### v1

@@ III Stationary States of Graphene Nanobubbles
188 1 0 1.476 3.57 31.54 3.52 0.113
2 0.0278 1.492 6.87 28.87 3.28 0.231
260 1 0 1.361 3.58 36.85 3.60 0.097
2 0.0216 1.367 7.15 32.62 3.34 0.219
398 1 0 1.296 3.58 45.58 3.66 0.079
2 0.0061 1.096 7.21 38.37 3.45 0.188
1 0 1.227 3.59 56.15 3.75 0.064
614 2 -0.0009 1.133 7.23 45.86 3.50 0.158
3 0.0108 1.165 10.36 44.91 3.41 0.213
1 0 1.180 3.58 70.11 3.80 0.051
962 2 -0.0068 1.083 7.23 56.27 3.55 0.128
3 -0.0013 1.015 10.76 53.26 3.45 0.202
1 0 1.146 3.58 82.44 3.84 0.043
1335 2 -0.0117 1.006 7.24 65.08 3.57 0.111
3 -0.0104 0.965 10.88 60.08 3.49 0.181
4 00.0012 0.980 14.04 58.02 3.45 0.242
1 0 1.125 3.59 96.27 3.86 0.037
1828 2 -0.0160 0.953 7.25 74.47 3.62 0.097
3 -0.0306 0.880 10.93 67.94 3.51 0.161
4 -0.0124 0.925 14.36 65.59 3.48 0.219
1 0 1.090 3.60 120.99 3.90 0.030
2 -0.0198 0.897 7.27 92.58 3.65 0.079
2909 3 -0.0238 0.860 10.89 83.19 3.55 0.131
4 -0.0228 0.841 14.55 79.27 3.50 0.184
5 -0.0171 0.846 17.81 76.76 3.47 0.232
1 0 1.077 3.60 140.68 3.91 0.026
2 -0.0218 0.872 7.26 106.36 3.67 0.068
3936 3 -0.0272 0.828 10.88 95.06 3.56 0.114
4 -0.0286 0.791 14.53 87.11 3.52 0.167
5 -0.0255 0.794 17.79 88.19 3.47 0.202
6 -0.0233 0.804 20.81 85.72 3.48 0.243
@@
The maximum possible number of layers, $l_{m}$, depends on the number of internal atoms, $N_{g}$ ($l_{m}$ increases with increasing $N_{g}$). For argon atoms, only single-layer packings ($l_{m}=1$) are possible for $N_{g}\leq 150$, while for $150<N_{g}<900$, single- and double-layer packings ($l_{m}=2$) are possible. For $900<N_{g}<1800$, the maximum number of layers is $l_{m}=3$; for $1800<N_{g}<2800$, it is $l_{m}=4$; and for $2800<N_{g}<4000$, it is $l_{m}=5$. A single-layer packing exists for any number of atoms. Therefore, the relative energy of a stationary state can be measured from the energy of the single-layer packing: $\Delta E=(E_{l}-E_{1})/N_{g}$, where $E_{l}$ is the energy of the $l$-layer stationary state.
@@
The dependence of the relative specific energy $\Delta E$, internal pressure $P$, height $H$, radius $R$, and density $d$ of the nanobubble’s stationary state on the number of argon atoms $N_{g}$ is given in Table 3. As can be seen from the table, single-layer stationary states are the most energetically favorable for $N_{g}<900$. For $900<N_{g}<3000$, two-layer states become more favorable, and for $N_{g}>3000$, three- and four-layer states are more favorable. The pressure inside the encapsulation monotonically decreases with an increasing number of atoms. With an increasing number of layers, a slight decrease in pressure and density is observed. The height of the nanobubble, $H$, always takes a discrete set of values proportional to the number of layers $l$: $H\approx h_{0}+(l-1)\Delta h$.
@@
For krypton atoms, the dependence of $\Delta E$, $P$, $H$, $R$, and $d$ on the number of internal atoms $N_{g}$ is presented in Table 4. Here, for $N_{g}\leq 140$, the maximum number of layers is $l_{m}=1$; for $140<N_{g}<600$, $l_{m}=2$; for $600<N_{g}<1300$, $l_{m}=3$; for $1300<N_{g}<2800$, $l_{m}=4$; and for $2800<N_{g}<4000$, $l_{m}=5$. As can be seen from the table, single-layer stationary states are the most energetically favorable for $N_{g}<900$, while two-layer states are favorable for $900<N_{g}<4000$. Here too, the pressure inside the nanobubble monotonically decreases with an increasing number of atoms. Pressure and density reach their maximum values for single-layer states.
For xenon atoms, the dependence of $\Delta E$, $P$, $H$, $R$, and $d$ on $N_{g}$ is presented in Table 5. Here, for $N_{g}\leq 116$, the maximum number of layers is $l_{m}=1$; for $116<N_{g}<600$, $l_{m}=2$; for $600<N_{g}<1300$, $l_{m}=3$; for $1300<N_{g}<2800$, $l_{m}=4$; for $2800<N_{g}<3800$, $l_{m}=5$; and for $3800<N_{g}<4000$, $l_{m}=6$. As can be seen from the table, single-layer stationary states are the most energetically favorable for $N_{g}<600$, two-layer states are favorable for $600<N_{g}<1400$, and three-layer states are favorable thereafter. Here too, the pressure inside the nanobubble monotonically decreases with an increasing number of atoms, and pressure and density reach their maximum values for single-layer packings.
@@
Figure 6: Dependence on temperature $T$ of (a) height $H$, (b) density $d$, and (c) internal pressure $P$ for a graphene nanobubble with $N_{g}=2909$ xenon atoms. Curves 1, 6, 11 correspond to the state with $l=1$; curves 2, 7, 12 – to $l=2$, …, and curves 5, 10, 15 – to $l=5$ layers. The dashed vertical lines correspond to the transition temperatures between states $l\rightarrow 4$: $T_{1}=960$, $T_{2}=390$, $T_{3}=330$, and $T_{5}=210$ K. The dotted vertical lines correspond to the melting temperature of the single-layer (420 K) and four-layer state (270 K).
@@
Figure 7: Dependence on temperature $T$ of the vertical displacement distribution $p(z)$ for $l$-layer states of $N_{g}=2909$ internal xenon atoms: (a) $l=1$; (b) $l=2$; (c) $l=3$; and (d) $l=4$.
@@ IV Stability of Stationary States to Thermal Fluctuations
Figure 8: Temperature dependence $T$ of the transverse displacement distribution $p(r)$ from the center of mass for $l$-layer states of $N_{g}=2909$ inner xenon atoms: (a) $l=1$; (b) $l=2$; (c) $l=3$; and (d) $l=4$.
Figure 9: Temperature dependence $T$ of the transverse displacement distribution $p(r)$ from the center of mass for $l$-layer states of $N_{g}=2909$ inner argon atoms: (a) $l=1$; (b) $l=2$; (c) $l=3$; and (d) $l=4$.
Dynamics of a thermalized graphene sheet with encapsulated atoms is described by the system of Langevin equations:
$M_{n}\ddot{\bf u}_{n}=-\frac{\partial H}{\partial{\bf u}_{n}}-\Gamma M_{n}\dot{\bf u}_{n}+\Xi_{n},~~n=1,...,N,$ (13)
@@
We numerically integrate the system of equations of motion (13) with initial conditions corresponding to a stationary nanobubble state:
@@
The equations of motion Eq. (13) with initial conditions (14) are solved numerically using the velocity Verlet method Verlet1967 . A time step of 1 fs is used in the simulations, as further reduction of the time step has no appreciable effect on the results.
Numerical simulations of the dynamics showed that among the $l$-layer states of the nanobubble, there is always one ”ground” state, which smoothly transitions into a layerless liquid state as the temperature increases. All other states transition into this ground state at a specific temperature. The dependence of the nanobubble’s height $H$, density $d$, and internal pressure $P$ on temperature is shown in Fig. 6.
@@
In Tables 3, 4, 5, the rows corresponding to the ground states are given in italics. Note that the ground states are not always the most energetically favorable. Their key feature is the smooth (non-abrupt) change in their shape with increasing temperature.
@@
Radial displacements also show changes in the distribution shape at the transition temperature $T_{l}$. The transition to a state with a larger number of layers leads to a sharp decrease in the diameter of the inner atom cluster—see Figs. 8 (a, b, c) and 9 (a, b, c). For the ground state with $l=4$ layers, only a continuous, gradual increase in the cluster diameter (its thermal expansion) occurs—see Figs. 8 (d) and 9 (d).
@@
The shape of the radial distribution also allows one to judge the phase state of the cluster. At low temperature, the radial distribution has a discrete structure characteristic of a crystalline (solid) state. Upon reaching the melting temperature $T_{l,m}$, the discrete distribution shape is replaced by a continuous one, characteristic of a liquid state. For $N_{g}=2909$ xenon atoms, the melting of the single-layer cluster occurs at $T_{1,m}=420\div 450$ K, and of the four-layer cluster at $T_{4,m}=270\div 300$ K—see Fig. 8. For argon atoms, the melting temperatures are $T_{1,m}=240\div 270$ K, $T_{4,m}=150\div 180$ K (see Fig. 9), and for krypton $T_{1,m}=300\div 330$ K, $T_{4,m}=210\div 240$ K. Such high melting temperatures are explained by the high value of internal pressure—see Fig. 6 (c).
@@
Figure 10: Temperature dependence $T$ of the ratio $H/R$ for the ground state of the nanobubble: with $N_{g}=962$ and 2909 argon atoms (curves 1 and 2); with $N_{g}=3936$ krypton atoms (curve 3); with $N_{g}=2909$ xenon atoms (curve 4).
@@
The existence of several stable states with different numbers of atom layers $l$ at low temperatures indicates that the nanobubbles here cannot have a universal (scalable) shape, and the height-to-radius ratio is not constant. Indeed, for a single-layer state ($l=1$), increasing the number of internal atoms $N_{g}$ does not change the nanobubble’s height ($H\approx 3.3$ Å, while its radius $R$ grows as $\sqrt{N_{g}}$; therefore, the ratio $H/R\rightarrow 0$ as $N_{g}\rightarrow\infty$. Based on the data in Tables 3, 4, and 5, the ratio $H/R$ can vary from 0 to 0.24 here. The violation of the scaling rule for graphene bubbles with a radius of $\sim 1$ nm was noted in the experimental study Villarreal2021 , where they were created via the implantation of noble gas ions He, Ne, and Ar. As can be seen from the table data, the relation $H/R\approx 0.2$ holds only for the ground states of the nanobubble.
@@
The dependence of the ratio $H/R$ on temperature $T$ for the ground states of the nanobubble is shown in Fig. 10. As can be seen from the figure, the value of $H/R$ increases monotonically with temperature, while the spread of values decreases: $H/R$ varies within the intervals [0.163, 0.206] at $T=0$ K; [0.192, 0.208] at $T=300$ K; [0.211, 0.217] at $T=600$ K. These estimates are in good agreement with the universal value $H/R=0.204$ obtained in Jain2017 for a graphene nanobubble with an ideal gas. Note that the study Korneva2023 reported an estimate of $H/R\approx 0.13$ for graphene nanobubbles with argon atoms at $T=2$ K, indicating that the simulation used not a ground stationary state of the nanobubble, but rather a state with a smaller number of layers.

### v3

@@ III Stationary States of Graphene Nanobubbles
125 1 0 1.545 4.01 25.39 2.99 0.158
2 0.0347 1.646 6.69 24.09 2.85 0.278
250 1 0 1.324 4.01 35.82 3.14 0.112
2 0.0134 1.411 7.54 31.48 3.04 0.240
500 1 0 1.182 4.02 50.36 3.27 0.080
2 -0.0081 1.238 7.64 41.87 3.19 0.182
3 0.0087 1.265 10.51 40.62 3.13 0.259
1 0 1.087 4.02 71.04 3.36 0.057
1000 2 -0.0226 1.113 7.64 56.73 3.30 0.135
3 -0.0154 1.141 11.23 52.79 3.25 0.213
1 0 1.021 4.02 100.34 3.42 0.040
2000 2 -0.0298 1.012 7.64 77.41 3.38 0.099
3 -0.0322 1.047 11.27 70.32 3.34 0.160
4 -0.0269 1.060 14.74 68.33 3.30 0.216
1 0 0.976 4.02 141.59 3.48 0.028
2 -0.0350 0.948 7.65 107.08 3.44 0.071
4000 3 -0.0411 0.977 11.28 96.85 3.38 0.116
4 -0.0412 0.989 14.94 91.93 3.35 0.163
5 -0.0380 0.995 18.43 87.79 3.34 0.210
6 -0.0345 0.996 20.81 86.63 3.33 0.240
Table 8: Dependence on the adhesion energy $\varepsilon_{0}$ and the number of layers $l$ of the relative specific energy $\Delta E$, pressure $P$, height $H$, radius $R$, and density $d$ of the stationary state of the nanobubble with $N_{g}=4000$ argon atoms.
$\varepsilon_{0}$ (eV)   $l$ $\Delta E$ (eV) $P$ (GPa) $H$ (Å) $R$ (Å) $d$ (g/cm3) $H/R$
1 0 0.684 3.27 119.76 1.81 0.027
2 0.0165 0.709 6.27 92.22 1.78 0.068
0.037 3 0.0035 0.729 9.27 83.28 1.76 0.111
4 0.0061 0.728 12.30 79.03 1.75 0.156
5 0.0095 0.727 15.00 75.83 1.74 0.198
1 0 1.407 3.17 112.57 2.12 0.028
2 -0.0247 1.118 6.14 85.36 2.05 0.072
0.148 3 -0.0302 1.209 9.14 77.63 2.00 0.118
4 -0.0317 1.241 12.18 73.43 1.96 0.166
5 -0.0310 1.249 15.08 69.62 1.95 0.217
The maximum possible number of layers, $l_{m}$, depends on the number of encapsulated atoms, $N_{g}$, increasing monotonically with $N_{g}$.
For helium atoms, the following behavior is observed: for $N_{g}<N_{1}=188$, only single-layer packings are possible ($l_{m}=1$); for $N_{1}<N_{g}<N_{2}=850$, both single- and double-layer packings are possible ($l_{m}=2$); for $N_{2}<N_{g}<N_{3}=2500$, the maximum number of layers is $l_{m}=3$; for $N_{3}<N_{g}<N_{4}=3500$, it is $l_{m}=4$; and for $N_{4}<N_{g}<4000$, it is $l_{m}=5$. The boundary values of the number of atoms, $N_{k}$ for the other noble gas are as follows: $N_{k}=182$, 650, 1500, 3500 for Ne; $N_{k}=145$, 550, 1600, 2500 for Ar; $N_{k}=123$, 450, 1300, 2400 for Kr; and $N_{k}=113$, 430, 1200, 2200 for Xe.
@@
A single-layer packing exists for any number of atoms. Therefore, the relative energy of a stationary state can be measured from the energy of the single-layer packing: $\Delta E=(E_{l}-E_{1})/N_{g}$, where $E_{l}$ is the energy of the $l$-layer stationary state.
@@
The dependence of the relative specific energy $\Delta E$, internal pressure $P$, height $H$, radius $R$, and density $d$ of the nanobubble’s stationary state on the number of helium atoms $N_{g}$ is given in Table 3. As can be seen from the table, single-layer stationary states are the most energetically favorable for $N_{g}\leq 2000$. The pressure inside the encapsulation monotonically decreases with an increasing number of atoms. With an increasing number of layers, a slight decrease in pressure and density is observed. The height of the nanobubble, $H$, always takes a discrete set of values proportional to the number of layers $l$: $H\approx l\Delta h$.
@@
For neon, argon, krypton and xenon atoms, the dependencies of $\Delta E$, $P$, $H$, $R$, and $d$ on the number of encapsulated atoms $N_{g}$ are presented in Tables 4, 5, 6 and 7, respectively. As can be seen from the tables, for all encapsulated atoms, the pressure inside the nanobubble decreases monotonically with an increasing number of atoms. Pressure and density reach their maximum values for single-layer states.
For neon atoms, single-layer stationary states are the most energetically favorable across the entire range investigated ($N_{g}\leq 4000$). For argon atoms, single-layer stationary states are the most energetically favorable only for $N_{g}<800$. For $800<N_{g}\leq 3000$, two-layer states become more favorable, and for $N_{g}>3000$, three-layer states are preferred. For krypton atoms, single-layer stationary states are the most energetically favorable for $N_{g}<700$. Two-layer states are favorable for $700<N_{g}<3000$, and three-layer states are favored for $N_{g}>3000$. For xenon atoms, single-layer stationary states are the most energetically favorable for $N_{g}<500$, while two-layer states are favorable for $500<N_{g}<2000$. For $N_{g}>2000$, three- and four-layer states become increasingly favorable.
@@
Figure 6: Dependence on temperature $T$ of (a) height $H$, (b) density $d$, and (c) internal pressure $P$ for a graphene nanobubble with $N_{g}=4000$ argon atoms. Curves 1, 6, 11 correspond to the state with $l=1$; curves 2, 7, 12 – to $l=2$, …, and curves 5, 10, 15 – to $l=5$ layers. The dashed vertical lines correspond to the transition temperatures between states $l\rightarrow 4$: $T_{1}=630$, $T_{2}=270$, $T_{3}=180$, and $T_{5}=120$ K. The dotted vertical lines correspond to the melting temperature of the single-layer ($T_{1,m}=255$ K) and four-layer state ($T_{4,m}=160$ K).
Figure 7: Dependence on temperature $T$ of height of graphene nanobubble $H$ with $N_{g}=4000$ internal (a) helium, (b) neon, (c) krypton and (d) xenon atoms (curves 1,…, 5 correspond to the state with l=1l=1,…, 5 layers.
@@
Figure 8: Dependence on temperature $T$ of the vertical displacement distribution $p(z)$ for $l$-layer states of $N_{g}=4000$ internal argon atoms: (a) $l=1$; (b) $l=2$; (c) $l=3$; and (d) $l=4$.
To understand how the shape of a nanobubble’s stationary state depends on the value of the interaction energy between the graphene sheet and the substrate, we also consider states with twice weaker and twice stronger interactions: at $\epsilon_{0}=0.037$ and $0.148$ eV. The weak interaction is characteristic of substrates with low density. For example, for an Ih ice substrate, the energy is $\epsilon_{0}=0.029$ eV Savin2019 . The strongest interaction occurs with the surface of a nickel crystal [with Ni(111) surface], where $\varepsilon_{0}=0.133$ eV Lahiri2011 . Thus, the three values of adhesion energy $\varepsilon=0.037$, 0.074, and 0.148 eV correspond to weak, medium (the most typical), and strong interaction of the graphene sheet with the substrate.
The numerical solution of the problem (12) showed that for all three characteristic values of the adhesion energy, a nanobubble containing $N_{g}=4000$ argon atoms can only exist in stationary states with l=1,,5l=1,...,5 layers; i.e., the maximum possible number of layers is $l_{m}=5$. The dependence of $\Delta E$, $P$, $H$, $R$, and $d$ on the number of layers is presented in Tables 5 and 8. As can be seen from the tables, the shape of the nanobubble’s stationary state depends only weakly on the value of $\epsilon_{0}$. A twofold increase in the adhesion energy leads to an increase in the internal pressure $P$, accompanied by a slight decrease in the nanobubble radius $R$. With weak adhesion, the most energetically favorable state is a single-layer nanobubble; with medium adhesion, a three-layer state; and with strong adhesion, a five-layer stationary state is the most favorable. Thus, an increase in the interaction energy between the graphene sheet and the substrate does not significantly change the step-like shape of the stationary states, but it noticeably increases the internal pressure and makes multilayer states energetically more favorable.
@@ IV Stability of Stationary States to Thermal Fluctuations
Figure 9: Temperature dependence $T$ of the transverse displacement distribution $p(r)$ from the center of mass for $l$-layer states of $N_{g}=4000$ inner argon atoms: (a) $l=1$; (b) $l=2$; (c) $l=3$; and (d) $l=4$.
Let us consider the dynamics of a molecular system in which a graphene sheet interacts with a Langevin thermostat. Dynamics of a thermalized graphene sheet with encapsulated atoms is described by the system of Langevin equations:
$M_{n}\ddot{\bf u}_{n}$ $=$ $-\frac{\partial H}{\partial{\bf u}_{n}}-\Gamma M_{n}\dot{\bf u}_{n}+\Xi_{n},~~n=1,...,N_{c},$~~~ (13)
$M_{n}\ddot{\bf u}_{n}$ $=$ $-\frac{\partial H}{\partial{\bf u}_{n}},~~n=N_{c}+1,...,N,$ (14)
@@
We numerically integrate the system of equations of motion (13), (14) with initial conditions corresponding to a stationary nanobubble state:
@@
The equations of motion Eq. (13), (14) with initial conditions (15) are solved numerically using the velocity Verlet method Verlet1967 . A time step of 1 fs is used in the simulations, as further reduction of the time step has no appreciable effect on the results.
Numerical simulations of the dynamics showed that among the $l$-layer states of the nanobubble, there is always one ”ground” state, which smoothly transitions into a layerless liquid configuration as the temperature increases. All other states transition into this ground state at a characteristic temperature $T_{l}$. The dependence of the nanobubble’s height $H$, density $d$, and internal pressure $P$ on temperature is shown in Fig. 6.
As can be seen from the figure, for a bubble with $N_{g}=4000$ argon atoms, a single-layer packing ($l=1$) persists up to a very high temperature $T_{1}=630$ K. Further temperature increase leads to its transition into a layerless molten configuration, accompanied by a sharp increase in height $H$, a decrease in density $d$, and internal pressure $P$. Two-layer atom packings ($l=2$) persist up to a temperature $T_{2}=270$ K, at which they transition into a four-layer packing. This transition is also accompanied by a sharp increase in height and a decrease in density. Three-layer packings ($l=3$) persist up to a temperature of $T_{3}=180$ K, and five-layer packings ($l=5$) persist up to $T_{5}=120$ K (see Fig. 6 (a)). Here, the ground state is the four-layer $(l=4)$ packing. Note that with weak adhesion $\varepsilon_{0}=0.037$ eV, the ground state is the three-layer ($l=3$) state, while with strong adhesion $\varepsilon_{0}=0.148$ eV, it is the five-layer ($l=5$) state.
@@
In Tables 3 to 8, the number of layers $l$ corresponding to the ground states is highlighted in bold. Note that the ground states are not always the most energetically favorable. Their key feature is the smooth (non-abrupt) change in their shape with increasing temperature. The ground state can also be obtained from the layerless configuration of the nanobubble at $T=900$ K by gradually reducing the thermostat temperature to zero.
@@
Radial displacements also show changes in the distribution shape at the transition temperature $T_{l}$. The transition to a state with a larger number of layers leads to a sharp decrease in the diameter of the inner atom cluster—see Fig. 9 (a,b,c). For the ground state with $l=4$ layers, only a continuous, gradual increase in the cluster diameter (its thermal expansion) occurs—see Fig. 9 (d).
@@
The shape of the radial distribution also allows one to judge the phase state of the cluster. At low temperature, the radial distribution has a discrete structure characteristic of a crystalline (solid) state. Upon reaching the melting temperature $T_{l,m}$, the discrete distribution shape is replaced by a continuous one, characteristic of a liquid state. For $N_{g}=4000$ argon atoms, the melting of the single-layer cluster occurs at $T_{1,m}\in$(240,270), and of the four-layer cluster at $T_{4,m}\in(150,180)$ K—see Fig. 9. For helium atoms, the melting temperatures are $T_{1,m}\approx 80$, $T_{4,m}\approx 48$, for neon $T_{1,m}\approx 135$, $T_{4,m}\approx 75$, for krypton $T_{1,m}\approx 315$, $T_{4,m}\approx 240$, and for xenon $T_{1,m}\approx 435$, $T_{4,m}\approx 285$ K.
@@
Figure 10: Temperature dependence $T$ of the ratio $H/R$ for the ground state of the nanobubble with (a) $N_{g}=500$, 1000, 2000, 4000 argon atoms (curves 1, 2, 3, 4) and (b) with $N_{g}=4000$ helium, neon, krypton, xenon atoms (curve 5, 6, 7, 8).
Figure 11: Temperature dependence $T$ of the ratio $H/R$ for the ground state of the nanobubble with $N_{g}=4000$ argon atoms for adhesion $\epsilon_{0}=0.037$, 0.074, 0.148 eV (curves 1, 2, 3)
The obtained values of the one-layer stability temperature and the ground state melting temperature do not depend on the number of atoms in the encapsulated cluster; the same values are obtained for $N_{g}=500$, 1000, 2000, and 3000. Therefore, it can be expected that the obtained estimates will also hold for larger clusters (for $N_{g}>4000$). Noble gases transition from the solid to the liquid state at very low temperatures (24, 84, 114, and 161 K for Ne, Ar, Kr, and Xe, respectively). The high melting temperatures of their encapsulated clusters are explained by the presence of high internal pressure and the effect of strong spatial confinement Faraji2022 .
The performed simulations show that the encapsulated noble gas cluster (graphene nanobubble) ceases to be a multistable system only at high temperatures $T>T_{1}$. Thus, at room temperature ($T=300$ K), graphene nanobubbles of Ne, Ar, Kr, and Xe, in addition to the ground three-dimensional state, also have a stable quasi-two-dimensional one-layer state—see Figs. 6 (a) and 7 (b, c, d).
@@
The existence of several stable states with different numbers of atomic layers $l$ at low temperatures indicates that nanobubbles in this regime cannot possess a universal (scalable) shape; consequently, the height-to-radius ratio is not constant. Indeed, for a single-layer state ($l=1$), increasing the number of encapsulated atoms $N_{g}$ leaves the nanobubble’s height unchanged ($H=2.2$, 2.5, 3.2, 3.5, and 4.0 Å for He, Ne, Ar, Kr, and Xe, respectively), while its radius $R$ grows as $\sqrt{N_{g}}$. The ratio $H/R$ therefore approaches zero as $N_{g}\rightarrow\infty$.
Based on the data in Tables 3 to 8, the ratio $H/R$ can vary from 0 to 0.28 here. The violation of the scaling rule for graphene bubbles with a radius of $\sim 1$ nm was noted in the experimental study Villarreal2021 , where they were created via the implantation of noble gas ions (He, Ne, and Ar). As can be seen from the table data, the relation $H/R\approx 0.2$ holds only for the ground states of the nanobubble.
@@
The dependence of the ratio $H/R$ on temperature $T$ for the ground states of the nanobubble is shown in Fig. 10. As can be seen from the figure, the value of $H/R$ increases monotonically with temperature, while the spread of values decreases: $H/R$ varies within the intervals [0.151, 0.206] at $T=0$; [0.174, 0.212] at $T=300$; [0.204, 0.220] at $T=600$; and [0.212, 0.222] at $T=900$ K. These estimates are in good agreement with the universal value $H/R=0.204$ obtained in Jain2017 for a graphene nanobubble containing an ideal gas. Note that the study Korneva2023 reported an estimate of $H/R\approx 0.13$ for graphene nanobubbles with argon atoms at $T=2$ K, indicating that the simulation used not the ground stationary state of the nanobubble, but rather a state with a smaller number of layers.
Note that the shape of the nanobubble’s ground state depends on the interaction energy $\epsilon_{0}$ between the graphene sheet and the substrate. The temperature dependence of the $H/R$ ratio for three characteristic values of the adhesion energy $\epsilon_{0}=0.037$, $0.074$, $0.148$ eV is shown in Fig. 11. As can be seen from the figure, an increase in the adhesion energy leads to a monotonic increase in the $H/R$ ratio. For example, at a temperature of $T=300$ K, the $H/R$ ratio can increase from $0.163$ for weak adhesion to $0.222$ for strong adhesion; at $T=600$ K — from $0.177$ to $0.238$; and at $T=900$ K — from $0.182$ to $0.247$. This monotonic dependence of the $H/R$ ratio on $\epsilon_{0}$ allows for experimental estimates of the adhesion between the graphene sheet and the substrate Wang2016 .